OpenStudy (anonymous):

Which equation represents the oxidation half-reaction for this redox reaction? Mg + CuSO4 d MgSOÁ + Cu -Mg ->Mg -Mg -> Mg+O4 -Mg -> Mg+2e- -Mg -> Mg+3e-

OpenStudy (anonymous):

*Mg+CuSo4 -> Mg SO4+Cu is what the equation should say..

OpenStudy (aaronq):

Do you know how to assign oxidation states?

OpenStudy (abb0t):

no. can you teach me?

OpenStudy (anonymous):

@aaronq Not really. I went through a tutorial on my online school and I'm still not grasping the concept as I wish I could.

OpenStudy (aaronq):

okay, i wrote something the other day for someone else, let me see if i can find it.

OpenStudy (aaronq):

You start off with a basic set of rules and you figure the rest out with simple arithmetic. Rules (most often true, except in rare occasions): Group 1 always +1 Group 2 always +2 Oxygen always -2 Group 7 always -1 Noble gases, elemental or homonuclear (the same element) molecules are zero The rest you have to figure out by subtracting or adding to achieve whatever charge the compound bears. eg. H2O H is group 1, so H = +1, there are 2 of them O must balance the charge so, O=-2 HCl H=+1, Cl=-1 MgSO4, you have to know that SO4^2- is a polyatomic ion, treat that as it's own separate thing: O=-2 Total contribution of O=-2(4) = -8 ^4 oxygen atoms Charge on ion=-2, so S has to fulfil the rest which is 6, so S=+6 back to MgSO4, so we treat SO4 as entity and it bears a -2 charge, Mg therefore has to be +2 to balance the charges.

OpenStudy (anonymous):

Oh okay~

OpenStudy (aaronq):

lol you'll be able to do it with some practice

OpenStudy (anonymous):

Thanks a lot~ I applied what you told me and got the answer right~ :)

OpenStudy (aaronq):

sweet ! :)